In a baby bouncer door swing, the child bounces in a harness suspended from a doorframe by a spring constant.

(a) If the spring stretches 0.242 m while supporting an 8.1-kg child, what is its spring constant?

            (b) What is the time for one complete bounce of this child?

(c) What is the child’s maximum velocity if the amplitude of her bounce is 0.222 m?
Answer #1

a)kx=8.1g

k*.242=8.1*10

k=334.71N/m

b)w=(k/m)^.5

=(334.71/8.1)^.5

=6.43

time=2pi/w

=0.98s

c)maxm velocity=Aw

=.222*6.43

=1.43m/s